# A Survey on Statistical Theory of Deep Learning: Approximation, Training Dynamics, and Generative Models

### v1

@@ Abstract
In this article, we review the literature on statistical theories of neural networks from three perspectives. In the first part, results on excess risks for neural networks are reviewed in the nonparametric framework of regression or classification. These results rely on explicit constructions of neural networks, leading to fast convergence rates of excess risks, in that tools from the approximation theory are adopted. Through these constructions, the width and depth of the networks can be expressed in terms of sample size, data dimension, and function smoothness. Nonetheless, their underlying analysis only applies to the global minimizer in the highly non-convex landscape of deep neural networks. This motivates us to review the training dynamics of neural networks in the second part. Specifically, we review papers that attempt to answer “how the neural network trained via gradient-based methods finds the solution that can generalize well on unseen data.” In particular, two well-known paradigms are reviewed: the Neural Tangent Kernel (NTK) paradigm, and Mean-Field (MF) paradigm. In the last part, we review the most recent theoretical advancements in generative models including Generative Adversarial Networks (GANs), diffusion models, and in-context learning (ICL) in the Large Language Models (LLMs). The former two models are known to be the main pillars of the modern generative AI era, while ICL is a strong capability of LLMs in learning from a few examples in the context. Finally, we conclude the paper by suggesting several promising directions for deep learning theory.
@@ 1 Introduction
In recent years, the field of deep learning (Goodfellow et al., 2016), a subdomain of machine learning, has experienced a substantial evolution. Its impact has transcended traditional boundaries, leading to significant advancements in sectors such as healthcare (Esteva et al., 2019), finance (Heaton et al., 2017), autonomous systems (Grigorescu et al., 2020), and natural language processing (Otter et al., 2020). Neural networks, the mathematical abstractions of our brain, lie at the core of this progression. Nevertheless, amid the ongoing renaissance of artificial intelligence, neural networks have acquired an almost mythical status, spreading the misconception that they are more art than science. It is important to dispel this notion. While the applications of neural networks may evoke awe, they are firmly rooted in mathematical principles. In this context, the importance of deep learning theory becomes evident. Several key points underscore its significance.
@@ 1.1 Why theory is important?
    Deep learning is a dynamic and rapidly evolving field, producing hundreds of thousands of publications online. Today’s models are characterized by highly intricate network architectures comprising numerous complex sub-components. Amidst this complexity, it becomes crucial to comprehend the fundamental principles underlying these models. To achieve this understanding, placing these models within a unified mathematical framework is essential. Such a framework serves as a valuable tool for distilling the core concepts from these intricate models, making it possible to extract and comprehend the key principles that drive their functionality.
@@
    Hyperparameters, such as learning rate, weight initializations, network architecture choices, activation functions, and batch sizes, significantly influence the quality of the estimated model. Understanding the proper ranges for these hyperparameters is essential, not only for theorists but also for practitioners. For instance, in the era of big data where there are millions of samples in one dataset, the theoretical wisdom gives us the depth of the network should scale logarithmically in sample size for the good estimation of compositional functions (Schmidt-Hieber, 2020).
In this review, we provide an overview of papers that delve into these concepts within precise mathematical settings, offering readers specific insights into the topics discussed above. Here, we try to avoid too many technicalities and make the introductions accessible to as many statisticians in various fields as possible.
@@ 1.2 Roadmap of the paper
We classify the existing literature on statistical theories for neural networks into three categories.
@@
    Approximation theory viewpoint. Recently, there have been a huge collection of works which bridge the approximation theory of neural network models (Yarotsky, 2017; Mhaskar, 1996; Petersen and Voigtlaender, 2018; Schmidt-Hieber, 2020; Montanelli and Du, 2019; Blanchard and Bennouna, 2022; Hornik et al., 1989; Hanin, 2019) with the tools in empirical processes (Van de Geer, 2000) to get the fast convergence rates of excess risks both in regression (Schmidt-Hieber, 2020; Hu et al., 2021) and classification (Hu et al., 2020; Kim et al., 2021) tasks under non-parametric settings. Approximation theory provides useful perspectives in measuring the fundamental complexities of neural networks for approximating functions in certain classes. Specifically, it enables the explicit constructions of neural networks for the function approximations so that we know how the network width, depth, and number of active parameters should scale in terms of sample size, data dimension, and the function smoothness index to get good convergence rates. For simplicity, we mainly consider the works in which the fully connected neural networks are used as the function estimators. These works include Schmidt-Hieber (2020); Kim et al. (2021); Shen et al. (2021); Jiao et al. (2021); Lu et al. (2021); Imaizumi and Fukumizu (2019, 2022); Suzuki (2018); Chen et al. (2019b); Suzuki and Nitanda (2021); Suh et al. (2022) under various problem settings. Yet, these works assume that the global minimizers of loss functions are obtainable, and are mainly interested in the statistical properties of these minimizers without any optimization concerns. However, this is a strong assumption, given the non-convexity of loss functions arising from the non-linearities of activation functions in the hidden layers.
@@
    Training dynamics viewpoint. Under this context, understanding the landscape of non-convex loss functions for neural network models and its implications on the generalization capabilities of neural networks becomes the next step in the literature. For example, a seminal empirical finding (Zhang et al., 2021) reveals that sufficiently overparameterized neural network models trained by stochastic gradient descent can fit the (noisy) data or even random noise perfectly but at the same time they can still generalize well. Among the many important discoveries on the role of overparameterizations such as Arora et al. (2019b); Jacot et al. (2018), revealed that the dynamics of deep neural networks with large enough width, trained via gradient descent (GD) in $\ell_{2}$-loss, behave similarly to those of functions in reproducing kernel Hilbert spaces (RKHS), where the kernel is associated with a specific network architecture. Many subsequent works study the training dynamics and the generalization abilities of neural networks in the kernel regime under various settings (Suh et al., 2021; Hu et al., 2021; Nitanda and Suzuki, 2020).
    Nonetheless, neural networks have been showing something more than just kernel regression, such as feature learning (Yang and Hu, 2020). This capability is achievable by permitting the network parameters to deviate significantly from their initial values, a freedom not afforded by the kernel regime. Many researchers try to fill in this gap (Ghorbani et al., 2020b; Wei et al., 2019), proving the statistical superiorities of neural networks over networks in kernel regime, yet still restricting the possible travelable distances of network parameters.
    Another important line of work tries to explain the learning dynamics of neural nets in a different regime, referred to as the Mean-Field (MF) regime. In this regime, the network parameters have the flexibility to deviate significantly from their initial values, even though it necessitates an infinite width. Finally, we conclude this section by introducing an unifying framework Yang and Hu (2020) for the kernel and MF regime, which gives us a comprehensive understanding of how the choices of weight initializations and scalings of learning rates in gradient-based methods can affect the dynamics of neural networks in its infinite width limit.
@@
    Generative modeling. In this section, we review the most recent theoretical advancements in generative models including Generative Adversarial Networks (GANs), Diffusion Models, and In-context learning in Large Language Models (LLMs). Over the past decade, GANs (Goodfellow et al., 2014) have stood out as a significant unsupervised learning approach, known for their ability to learn the data distributions and efficiently sample the data from it. In this review, we will introduce papers that studied statistical properties of GANs (Arora et al., 2017; Liang, 2021; Chen et al., 2020a; Bai et al., 2018; Zhang et al., 2017; Schreuder et al., 2021). Recently, another set of generative models, i.e., diffusion model, has shown superior performances in generating very good qualities of synthetic data in various data modalities including image (Song et al., 2020; Dhariwal and Nichol, 2021), tabular data (Kim et al., 2022; Suh et al., 2023), medical imaging (Müller-Franzes et al., 2022) etc., beating GAN based models by a large margin. However, given the model’s complex nature and its recent introduction in the community, the theoretical reason why it works so well remains vague. Lastly, we will review an interesting phenomenon commonly observed in Large Language Models referred to as In-context learning (ICL). It refers to the ability of LLMs conditioned on prompt sequence consisting of examples from a task (input-output pairs) along with the new query input, the LLM can generate the corresponding output accurately. Readers can refer to nice survey papers Gui et al. (2021); Yang et al. (2022) on the detailed descriptions of methodologies and the applications of GANs, and diffusion models in various domains. For an overview of ICL, see survey by Dong et al. (2022) which highlights some key findings and advancements in this direction.
@@ 1.3 Existing surveys on deep learning theory
To our knowledge, there are three existing survey papers (Bartlett et al., 2021; Fan et al., 2021; Belkin, 2021) on deep learning theory. There is an overlap in certain subjects covered by each of the papers, but their main focuses are different. Bartlett et al. (2021) provided a comprehensive and technical survey on statistical understandings of deep neural networks. In particular, the author’s focus was on examining the significant influence of over-parametrization in neural networks, which plays a key role in enabling gradient-based methods to discover interpolating solutions. These methods introduce implicit regularization, as discussed in Neyshabur (2017), or lead to a phenomenon known as benign overfitting Bartlett et al. (2020). Fan et al. (2021) introduced the most commonly employed neural network architectures in practice such as Convolutional Neural Net (CNN), Recurrent Neural Networks (RNN), and training techniques such as batch normalization, dropout from statistical viewpoints. A brief introduction to the approximation theory of neural networks is provided as well. Similarly as Bartlett et al. (2021)Belkin (2021) reviewed the role of overparametrizations for implicit regularization and benign overfitting observed not only in neural network models but also in classic statistical models such as weighted nearest neighbor predictors. Most notably, they provided intuitions on the roles of the overparametrizations of non-convex loss landscapes of neural networks through the lens of optimization.
2 Excess Risk Bounds for Neural Networks
We briefly outline fully connected networks, which are the main object of interest throughout this paper. From a high level, deep neural networks can be viewed as a family of nonlinear statistical models that can encode highly nontrivial representations of data. The specific network architecture $(L,\mathbf{p})$ consists of a positive integer $L$, called the number of hidden layers, and a width vector $\mathbf{p}:=(\mathbf{p}_{0},\dots,\mathbf{p}_{L+1})\in\mathbb{N}^{L+2}$. A fully connected neural network with architecture $\tilde{f}$ is then any function of the form for the input feature $\mathbf{x}\in\mathcal{X}$:
@@ 2 Excess Risk Bounds for Neural Networks
where $\mathbf{W}_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ is a weight matrix with $\mathbf{p}_{0}=d$, $\mathbf{p}_{L+1}=1$ and $\mathbf{v}_{i}\in\mathbb{R}^{p_{i}}$ is a shift vector.
Network functions are built by alternating matrix-vector multiplications with the action of the nonlinear activation function $\sigma$. Let $\|\mathbf{W}_{j}\|_{0}$ and $|\mathbf{v}_{j}|_{0}$ be the number of nonzero entries of $\mathbf{W}_{j}$ and $\mathbf{v}_{j}$ in the $j^{\text{th}}$ hidden layer. The final form of neural network we consider is given by:
@@
Here, the complexity of networks is mainly measured through the three metrics: $(1)$ the width denoted as $\mathbf{p}$, $(2)$ the depth, denoted as $L$, and $(3)$ the number of active parameters, denoted as $\mathcal{N}$ of the network. See Anthony and Bartlett (1999) for in-depth discussions on these metrics.
2.1 Preliminary Problem Setting

### v3

@@ Abstract
In this article, we review the literature on statistical theories of neural networks from three perspectives: approximation, training dynamics and generative models. In the first part, results on excess risks for neural networks are reviewed in the nonparametric framework of regression (and classification in Appendix B). These results rely on explicit constructions of neural networks, leading to fast convergence rates of excess risks. Nonetheless, their underlying analysis only applies to the global minimizer in the highly non-convex landscape of deep neural networks. This motivates us to review the training dynamics of neural networks in the second part. Specifically, we review papers that attempt to answer “how the neural network trained via gradient-based methods finds the solution that can generalize well on unseen data.” In particular, two well-known paradigms are reviewed: the Neural Tangent Kernel (NTK) paradigm, and Mean-Field (MF) paradigm. Last but not least, we review the most recent theoretical advancements in generative models including Generative Adversarial Networks (GANs), diffusion models, and in-context learning (ICL) in the Large Language Models (LLMs) from two perpsectives reviewed previously, i.e., approximation and training dynamics.
@@ 1 Introduction
In recent years, the field of deep learning (Goodfellow et al., 2016) has experienced a substantial evolution. Its impact has transcended traditional boundaries, leading to significant advancements in sectors such as healthcare (Esteva et al., 2019), finance (Heaton et al., 2017), autonomous systems (Grigorescu et al., 2020), and natural language processing (Otter et al., 2020). Neural networks, the mathematical abstractions of our brain, lie at the core of this progression. Nevertheless, amid the ongoing renaissance of artificial intelligence, neural networks have acquired an almost mythical status, spreading the misconception that they are more art than science. It is important to dispel this notion. While the applications of neural networks may evoke awe, they are firmly rooted in mathematical principles. In this context, the importance of deep learning theory becomes evident. Several key points underscore its significance.
@@ 1.1 Why theory is important?
    Deep learning is a dynamic and rapidly evolving field, producing hundreds of thousands of publications online. Today’s models are characterized by highly intricate network architectures comprising numerous complex sub-components (e.g., Transformer (Vaswani et al., 2017)). Amidst this complexity, it becomes crucial to comprehend the fundamental principles underlying these models. To achieve this understanding, placing these models within a unified mathematical framework is essential. Such a framework serves as a valuable tool for distilling the core concepts from these intricate models, making it possible to extract and comprehend the key principles that drive their functionality.
@@
    Hyperparameters, such as learning rate, weight initializations, network architecture choices, activation functions, etc, significantly influence the quality of the estimated model. Understanding the proper ranges for these hyperparameters is essential, not only for theorists but also for practitioners. For instance, in the era of big data where there are millions of samples in one dataset, the theoretical wisdom gives us the depth of the network should scale logarithmically in sample size for the good estimation of compositional functions (e.g., See Schmidt-Hieber (2020)).
In this review, we provide an overview of papers that delve into these concepts within mathematical settings, offering readers specific insights into the topics discussed above. Here, we try to avoid too many technicalities and make the introductions accessible to as many statisticians as possible. Some overly technical components are deferred to the Appendix.
@@ 1.2 Roadmap of the paper
We classify the existing literature on statistical theories of neural networks into three categories. (Section 2, Section 3, Section 4, respectively.)
@@
    Approximation theory viewpoint. Recently, there has been a huge collection of works which bridge the approximation theory of neural network models (Yarotsky, 2017; Mhaskar, 1996; Petersen and Voigtlaender, 2018; Schmidt-Hieber, 2020; Montanelli and Du, 2019; Blanchard and Bennouna, 2022; Hornik et al., 1989; Hanin, 2019) with the tools in empirical processes (Van de Geer, 2000) to obtain the fast convergence rates of excess risks both in regression (Schmidt-Hieber, 2020; Hu et al., 2021) and classification (Hu et al., 2020a; Kim et al., 2021) tasks under non-parametric settings. Approximation theory provides useful perspectives in measuring the fundamental complexities of neural networks for approximating functions in certain classes. Specifically, it enables the explicit constructions of neural networks for the function approximations so that we know how the network width, depth, and number of active parameters should scale in terms of sample size, data dimension, and the function smoothness index to get good statistical convergence rates. For simplicity, we mainly consider the works in which the fully connected neural networks are used as the function estimators. These works include Schmidt-Hieber (2020); Kim et al. (2021); Shen et al. (2021); Jiao et al. (2021); Lu et al. (2021); Imaizumi and Fukumizu (2019, 2022); Suzuki (2018); Chen et al. (2019); Suzuki and Nitanda (2021); Suh et al. (2022) under various problem settings. Yet, these works assume that the global minimizers of loss functions are obtainable, and are mainly interested in the statistical properties of these minimizers without any optimization concerns. However, this is a strong assumption, given the non-convexity of loss functions arising from the non-linearities of activation functions in the hidden layers.
@@
    Training dynamics viewpoint. Understanding the landscape of non-convex loss functions for neural network models and its impact on their generalization capabilities represents a critical next step in the literature. However, the non-trivial non-convexity of this landscape poses significant challenges for the mathematical analysis of many intriguing phenomena observed in neural networks. For example, a seminal empirical finding, Zhang et al. (2021), reveals that neural networks in their experiments trained on a standard image classification training set (CIFAR 10), can fit the (noisy) training data perfectly, but at the same time showing respectable prediction performance. (See Figure 1(c) in Zhang et al. (2021).) This contradicts the classic statistical wisdom of bias-variance tradeoff, which states the overfitted models cannot generalize well. The role of overparameterizations (e.g., Bartlett et al. (2021)) on the non-convex optimization landscape of neural networks has been intensively studied over the past years, and we review the relevant literatures under this context. For instance, Jacot et al. (2018) revealed that the dynamics of highly overparameterized neural networks with large enough width, trained via gradient descent (GD) in $\ell_{2}$-loss, behave similarly to those of functions in reproducing kernel Hilbert spaces (RKHS), where the kernel is associated with a specific network architecture. Many subsequent works study the training dynamics and the generalization abilities of neural networks in the kernel regime under various settings (Hu et al., 2021; Nitanda and Suzuki, 2020). However, due to technical constraints (as detailed in Subsection 3.1), networks in the kernel regime fail to explain the essential functionality of neural networks—feature learning (Zhong et al., 2016). Another important line of work focuses on understanding the learning dynamics of neural networks in the Mean-Field (MF) regime, where feature learning becomes more explainable. Nonetheless, the analysis within the MF regime is challenging to generalize to deep networks and requires infinite widths. Finally, we conclude this section by presenting several approaches that go beyond or unify the two regimes.
@@
    Generative modeling. In this section, we review the most recent theoretical advancements in generative models including Generative Adversarial Networks (GANs), Diffusion Models, and In-context learning in Large Language Models (LLMs). The works to be introduced are based on the philosophies of two paradigms (approximation & training dynamics). Over the past decade, GANs (Goodfellow et al., 2014) have stood out as a significant unsupervised learning approach, known for their ability to learn the data distributions and efficiently sample the data from it. In this review, we will introduce papers that studied statistical properties of GANs (Arora et al., 2017; Liang, 2021; Chen et al., 2020a; Bai et al., 2018; Zhang et al., 2017; Schreuder et al., 2021). Recently, another set of generative models, i.e., diffusion model, has shown superior performances in generating impressive qualities of synthetic data in various data modalities including image (Song et al., 2020; Dhariwal and Nichol, 2021), tabular data (Kim et al., 2022; Suh et al., 2023), medical imaging (Müller-Franzes et al., 2022) etc., beating GAN based models by a large margin. However, given the model’s complex nature and its recent introduction in the community, the theoretical reason why it works so well remains vague. Lastly, we will review an interesting phenomenon commonly observed in Large Language Models referred to as In-context learning (ICL). It refers to the ability of LLMs conditioned on prompt sequence consisting of examples from a task (input-output pairs) along with the new query input, the LLM can generate the corresponding output accurately. Readers can refer to nice survey papers Gui et al. (2021); Yang et al. (2022) on the detailed descriptions of methodologies and the applications of GANs, and diffusion models in various domains. For an overview of ICL, see survey by Dong et al. (2022) which highlights some key findings and advancements in this direction.
In relation to Subsection 1.1, the advantages of neural networks over classic statistical function estimators are primarily discussed in Sections 2 and 3 under various problem settings. In Section 3, we review the work of Yang and Hu (2020), which suggests appropriate parameter initialization scalings and learning rates for feature learning in large-scale (infinite width) neural networks.
@@ 1.3 Existing surveys on deep learning theory
To our knowledge, there are four existing survey papers (Bartlett et al., 2021; Fan et al., 2021; Belkin, 2021; He and Tao, 2020) on deep learning theory. There is an overlap in certain subjects covered by each of the papers, but their main focuses are different from each other. Bartlett et al. (2021) provided a comprehensive and technical survey on statistical understandings of deep neural networks. In particular, the authors focused on examining the significant influence of overparameterization in neural networks, which plays a key role in enabling gradient-based methods to discover interpolating solutions. Fan et al. (2021) introduced the most commonly employed neural network architectures in practice such as Convolutional Neural Net (CNN), Recurrent Neural Networks (RNN), and training techniques such as batch normalization, and dropout. A brief introduction to the approximation theory of neural networks is provided as well. Similarly as Bartlett et al. (2021)Belkin (2021) reviewed the role of overparameterizations for implicit regularization and benign overfitting observed not only in neural network models but also in classic statistical models such as weighted nearest neighbor predictors. Most notably, they provided intuitions on the roles of the overparameterization of non-convex loss landscapes of neural networks through the lens of optimization. He and Tao (2020) provides a comprehensive overview of deep learning theory, including the ethics and security problems that arise in data science and their relationships with deep learning theory. We recommend readers review all these papers to gain a comprehensive understanding of this emerging field. Our paper offers a unique and comprehensive survey of the statistical results of neural networks, focusing on approximation theory and training dynamics, while also covering generative models within these two paradigms.
2 Approximation theory-based statistical guarantees
We outline fully connected networks, which are the main object of interest throughout this review. From a high level, deep neural networks can be viewed as a family of nonlinear statistical models that can encode highly nontrivial representations of data. The specific network architecture $(L,\mathbf{p})$ consists of a positive integer $L$, called the number of hidden layers, and a width vector $\mathbf{p}:=(\mathbf{p}_{0},\dots,\mathbf{p}_{L+1})\in\mathbb{N}^{L+2}$, recording the number of nodes from input to output layers in the network. A fully connected neural network, $\tilde{f}$, is then any function of the form for the input feature $\mathbf{x}\in\mathcal{X}$:
@@ 2 Approximation theory-based statistical guarantees
where $\mathbf{W}_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ is a weight matrix with $\mathbf{p}_{0}=d$, $\mathbf{p}_{L+1}=1$ and $\mathbf{\sigma}$ is the non-linear activation function. Here, the activation function plays a key-role in neural network allowing the non-linear representations of the given data $\mathbf{x}$. Popular examples include: ReLU($x$)=$\max(x,0)$, Sigmoid($x$)=$\frac{1}{1+e^{-x}}$. We omit the bias terms added on the outputs of pre-activated hidden layers for simplicity. But bias terms are needed for universal approximation if the input data is not appended with a constant entry.
Under this setting, complexity of the networks is mainly measured through the three metrics: $(1)$ the maximum width, denoted as $\mathbf{p}_{\textbf{max}}:=\max_{i=0,\dots,L+1}\mathbf{p}_{i}$, $(2)$ the depth, denoted as $L$, and $(3)$ the number of non-zero parameters, denoted as $\mathcal{N}$. Letting $\|\mathbf{W}_{j}\|_{0}$ be the number of non-zero entries of $\mathbf{W}_{j}$ in the $j^{\text{th}}$ hidden layer, the final form of neural network we consider is given by:
@@
In the approximation theoretic literature, the capacity or expressive power of neural network is often characterized by the tuple $(L,\mathbf{p}_{\textbf{max}},\mathcal{N})$. Let $\mathcal{G}$ be a function class where the target function $f_{\star}$ belongs. The main question frequently asked is: given the fixed approximation error, $\varepsilon_{\text{Apprx}}$, defined as follows:
$\varepsilon_{\text{Apprx}}:=\sup_{f_{\star}\in\mathcal{G}}\inf_{f\in\mathcal{F}(L,\mathbf{p},\mathcal{N})}\left\|f-f_{\star}\right\|_{L_{p}},$ (3)
how does the network architecture $(L,\mathbf{p}_{\textbf{max}},\mathcal{N})$ scale in terms of $\varepsilon_{\text{Apprx}}$? (henceforth, the subscript Apprx will be omitted in $\varepsilon_{\text{Apprx}}$.) Note the sup is taken over the function class $\mathcal{G}$ and the inf is taken over the neural network class $\mathcal{F}$. The distance between two functions is measured via $L_{P}$-norm.
2.1 Expressive power of fully-connected networks
In this subsection, we briefly review some important results in the approximation theory of neural networks. For more comprehensive reviews on this topic, readers can refer to DeVore et al. (2021).
Approximating functions in $\mathcal{G}$: The specifications of function class $\mathcal{G}$ and $\mathcal{F}$ allow us to derive many interesting insights on the power of neural networks. For instance, the celebrated “universal approximation theorem” states that any continuous functions (i.e., $\mathcal{G}:=\{\text{Continuous Functions on $\mathbb{R}^{d}$}\}$) can be approximated by a shallow neural network (i.e., one-hidden layer) with sigmoid activation function (i.e., :=$\mathcal{F}:=\{\text{Shallow Neural Networks}\}$) at an arbitrary accuracy (Cybenko (1989); Hornik et al. (1989, 1990)). However, achieving a good approximation may require an extremely large number of hidden nodes, which significantly increases the capacity of $\mathcal{F}$.  Barron (1993, 1994) developed an approximation theory for function classes $\mathcal{G}$ with limited capacity, measured by the integrability of their Fourier transform. Interestingly, the approximation result is not affected by the dimension of input data $d$, and this observation matches with the experimental results that deep learning is very effective in dealing with high-dimensional data.
Nonetheless, the capacity of $\mathcal{G}$ in Barron (1994) is rather limited. Another typical route of the analysis is to specifying the smoothness of function classes. Readers can roughly understand that smoothness refers to the highest order of derivatives that the functions can possess. Notably, Yarotsky (2017) demonstrated that deep ReLU networks (1) cannot escape the curse of dimensionality when approximating functions in the unit ball in Sobolev space. They established that the order $\mathcal{N}=\mathcal{O}(\varepsilon^{-\frac{d}{r}})$ is sharp, with matching lower and upper bounds. Petersen and Voigtlaender (2018) generalized the results to the class of piece-wise smooth functions. Later, Schmidt-Hieber (2020) developed a theory for any network architecture satisfying the set of conditions on $(L,\mathbf{p}_{\textbf{max}},\mathcal{N})$, deep ReLU nets can achieve good approximation rates for functions in Hölder class. (More details on the technical results of Schmidt-Hieber (2020) are deferred in the Appendix A.) This should be contrasted to the Yarotsky (2017) where they proved the “existence” of a network with good approximation. Many researchers have been working on considering either more general (i.e., Besov space) or more specific (i.e., hierarchical compositional function) function classes $\mathcal{G}$ than Hölder class. These considerations have facilitated numerous intriguing comparisons between classic statistical function estimators and deep neural networks in terms of their fundamental limits, specifying the second item in Subsection 1.1.
The benefits of depth: It has been shown that the expressive power of deep neural networks grows with respect to the number of layers (i.e., $L$) by several studies. Delalleau and Bengio (2011) showed there exist families of functions that can be represented much more efficiently with a deep network than with a shallow one (i.e. with substantially fewer hidden units). In the asymptotic limit of depth, Pascanu et al. (2013) showed deep ReLU networks can represent exponentially more piece-wise linear functions than their single-layer counterparts can do, given that both networks have the same number of nodes. Later, Montufar et al. (2014) proved that a similar result can be derived with the fixed number of hidden layers. Poole et al. (2016) showed deep neural networks can disentangle highly curved manifolds in an input space into flat manifolds in a hidden space, while shallow networks cannot. Mhaskar et al. (2017) demonstrated that deep ReLU networks can approximate compositional functions with significantly fewer parameters (i.e., $\mathcal{N}$) than shallow neural networks need in order to achieve the same level of approximation accuracy.
Bounded-width: The effects of width on the expressive power of neural networks have recently been studied (Lu et al., 2017; Hanin, 2019; Kidger and Lyons, 2020; Park et al., 2020; Vardi et al., 2022).  Lu et al. (2017) showed that the minimal width for universal approximation (denoted as $w_{\text{min}}$) using ReLU networks w.r.t. the $L_{1}$ norm of functions from $\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $d+1\leq w_{\text{min}}\leq d+4$.  Kidger and Lyons (2020) extended the results to $L_{p}$-approximation of functions from $\mathbb{R}^{d}\rightarrow\mathbb{R}^{\text{out}}$, and obtained $w_{\text{min}}\leq d+d_{\text{out}}+1$.  Park et al. (2020) further improved $w_{\text{min}}=\max\{d+1,d_{\text{out}}\}$. Universal approximations of narrow networks with other activation functions had been studied in Park et al. (2020); Kidger and Lyons (2020); Johnson (2018). Note that the aforementioned works require the depth of networks to be exponential in input dimension with bounded-width, which are the dual-versions of universal approximation of bounded depth networks from Cybenko (1989); Hornik et al. (1989, 1990). Interestingly, Vardi et al. (2022) provided an evidence that width of the networks can be less important than depth for the expressive power of neural nets. They showed that the price for making the width small is only a linear increase in the network depth, in sharp contrast to the results mentioned earlier on how making the depth small may require an exponential increase in the network depth.
2.2 Statistical guarantees for regression tasks
The natural question is: what are the interpretations or consequences of the results in approximation theory for deriving statistical guarantees of neural networks under noisy observations? In this subsection, we focus on reviewing several important results under regression tasks in this regard. As preliminary, we introduce the settings frequently adopted in statistical learning theory.